OpenStudy (anonymous):

If it takes Sam two hours to clean a house and it takes Mike 5 hours to clean the same house, how long would it take if they worked together?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } + \frac{ 1 }{ 5 } = \frac{ 1 }{ x }\] Is this correct so far?

OpenStudy (anonymous):

that will be 7/10 =1/x so cross multiply gives you 7x=10 x=1 3/7

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Welcome and anytime!!!